OpenStudy (xmistermayhem):

A plane is descending towards a runway in Chicago. The plane is flying at an altitude of 11 miles and is 193 miles from the runway, as measured from the ground. What angle will the plane's path make with the runway?

OpenStudy (baru):



OpenStudy (xmistermayhem):

...? :s what?

OpenStudy (seascorpion1):

\[\tan(\theta)=\frac{opposite}{adjacent}\]hope this helps.

OpenStudy (michele_laino):

hint: we can write this equation: \[\huge 11 = 193 \cdot \tan \theta \]

OpenStudy (seascorpion1):

Can you solve it?

OpenStudy (seascorpion1):

@xmistermayhem

OpenStudy (xmistermayhem):

Uh. I got 0.0569948186528497

OpenStudy (seascorpion1):

tan(theta)=11/193 so theta=tan^-1(11/193)

OpenStudy (seascorpion1):

What have you got for theta?

OpenStudy (seascorpion1):

You seem to have worked out what tan theta is, do tan inverse of both sides.

OpenStudy (michele_laino):

please you have to do this computation: \[\huge \arctan \left( {0.056994818} \right)\]

OpenStudy (michele_laino):

or, using a different, and equivalent notation: \[\huge {\tan ^{ - 1}}\left( {0.056994818} \right)\]

OpenStudy (seascorpion1):

@xmistermayhem have you solved it yet?

OpenStudy (triciaal):

did you forget to close the question or do you need more help?